Check my work Required information [The following information applies to the questions displayed below.] Part 1 of 4 Vals HaI'm having an extremely hard time understanding income tax expense and income tax payable. Wouldn't you multiply net income by the tax rate percentage? I'm sure I'm missing a step somewhere. Can I get a step-by-step explanation on how to get the correct amount for income tax expense and income tax payable, please?
Answer #1
Income tax expense is calculated by multiplying income before tax by tax rate percentage
We need to calculate Income before income tax with all adjustments.
Service Revenues 64450
Less: Expenses
Salaries and wages expense 28500 =28400+100
Utilities expense 11880 =11500+380
Rent expense 15600 =1300*12
Supplies expense 6750 =4100+(3600-950)
Total expenses 62730
Income before income tax 1720
Income before income tax 1720
X Tax rate 25%
Income tax expense / payable 430